jhonyy9 (jhonyy9):

prove that for every natural numbers n greater than 2 allways will be (exist) two primes p and q so that this equation n=(p-1)/2 +(q-1)/2 +1 is true .

jhonyy9 (jhonyy9):

for example : 3=(3-1)/2 +(3-1)/2 +1 3=2/2 +2/2 +1 3=1+1+1 3=3

OpenStudy (anonymous):

I think this is related to Goldbach's conjecture that every even integer is the sum of two primes. http://en.wikipedia.org/wiki/Goldbach%27s_conjecture Since if \[ n = \frac {p-1} 2+ \frac {q-1}2 +1 \\ 2 n = p-1 + q-1 +2 =p+q \]

jhonyy9 (jhonyy9):

- so because every prims can be writing in the form of p=2x+1 and q=2y+1 where x and y are natural numbers so than : 2x+1-1 2y+1-1 n=-------- + -------- +1 2 2 n= 2x/2 +2y/2 +1 n=x+y+1 - so what can proving verx easy by induction that is true

jhonyy9 (jhonyy9):

very easy - sorry

OpenStudy (anonymous):

There is something wrong in your writing above, or I misunderstood your question. Your question and I am copying and pasting is "prove that for every natural numbers n greater than 2 allways will be (exist) two primes p and q so that this equation n=(p-1)/2 +(q-1)/2 +1 is true . " In your proof, you defined n the way you want to be, but the question was to prove that every n can be written as n=(p-1)/2 +(q-1)/2 +1 If you prove this, you would be very famous!!! You have to construct p and q from n and not vice versa.

jhonyy9 (jhonyy9):

- ok i understand it but you need to know again that n is an indifferent natural number greater than 2 - secondly by definition every primes greater than 2 can be writing in the form of 2n+1 ,where n is a natural number - 3. - by induction is very easy proving that every n ,natural numbers grater than 2 ,can be writing like a sum of two natural numbers x and y plus 1 , so like n=x+y+1, - 4. - so than p=2x+1 and q=2y+1 ,where p and q are primes grater than 2 and x,y are natural numbers ,so than the statement n = (p-1)/2 +(q-1)/2 +1 can be writing like n=(2x+1-1)/2 +(2y+1-1)/2 +1 n=2x/2 +2y/2 +1 n=x+y+1

OpenStudy (anonymous):

I said what I said before, It does not seem that you get it. Sorry.